If the mean score in a STAT 1 examination is 56 and the variance is 4. What is the coefficient of variation


1
Expert's answer
2021-11-25T11:29:50-0500
"CV= \dfrac{\sigma}{\mu} \cdot 100\%=\dfrac{\sqrt{4}}{56}\cdot 100\%"

"=\dfrac{25}{7}\%\approx3.5714\%"